OpenStudy (anonymous):

An artifact was found and tested for its carbon-14 content. If 82% of the original carbon-14 was still present, what is its probable age (to the nearest 100 years)? (Carbon-14 has a half-life of 5,730 years.)

OpenStudy (anonymous):

the equation for this is \[f=i(e ^{kt})\] First you need to find the rate of decay (k). Because half of it decays in 5730 years, I can say \[50=100 e^{k(5730)}\] so ln(1/2)=5730k k=-1.209*10^-4 Using this k, you can find the answer \[87=100e ^{-1.209*10^{-4}t}\] solving this gives \[ln(.87)={-1.209*10^{-4}t}\] t=1151.878 years